Problem 3 (10 points) Suppose a sequence satisfies the below given recurrence relation and initial conditions. Find an explicit formula for the sequence a -6a--9a,-2 for all integers k2 2 ao = 1, a1 = 3
Answer #1

Given Recurrence Relation is ak = 6 ak-1 - 9 ak-2  for k  \tiny \geq 2 and

initial conditions are a0 = 1 a1 = 3.

Here a2 = 6 a1 - 9 a0 = 6 * 3 - 9 * 1 = 18 - 9 = 9.

Here a3 = 6 a2 - 9 a1 = 6 * 9 - 9 * 3 = 54 - 27 = 27.

Here a4 = 6 a3 - 9 a2 = 6 * 27 - 9 * 9 = 162 - 81 = 81.

Here a0 = 1 = 30

a1 = 3 = 31

a2 = 9 = 32

a3 =  27 = 33

a4 = 81 = 34

In the same way an = 3n.

Hence explicit formula for the sequence is an = 3n.